OpenStudy (anonymous):

help me please! A total of 448 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (anonymous):

a + s = 448 s = 3a now sub 3a in for s in the first equation and solve for a

OpenStudy (anonymous):

I don't get it. so its 3a=448?

OpenStudy (anonymous):

a + s = 448 a + 3a = 448 ( a is the same as 1a....so add them) 4a = 448 (divide both sides by 4) a = 448/4 a = 112 do you understand what I did ?

OpenStudy (anonymous):

oh okay, yes I understand thank you!

OpenStudy (anonymous):

any time :)